OpenStudy (anonymous):

Find the solution set for y = 2x + 1, given the replacement set {(-2, -3), (-1, -1), (0, 2), (1, 3), (2, 6)}.

OpenStudy (anonymous):

1. just substitute the values of x from the replacement set to the equation y=2x+1 and solve for y.

OpenStudy (anonymous):

(-2,-3): y=2(-2)+1 which implies y=-4+1=-3 thus (-2,-3) is an element in your solution set. do this for the other ordered pairs